The coil is rotated through an angle of 90° in 0.2 second. Calculate the magnitude of the average e.m.f. induced in the coil while it is being rotated.


1
Expert's answer
2021-05-30T13:29:34-0400
"E=\frac{d\Phi}{dt}=\frac{\Delta\Phi}{\Delta t}=0.5\Delta\Phi"

Assume that "\Delta\Phi=8\ (Wb)" . Then

"E=0.5(8)=4\ V"